OpenStudy (anonymous):

There are five thieves, each loot a bakery one after the other such that the first one takes 1/2 of the total no. of the breads plus 1/2 of a bread. Similarly 2nd, 3rd,4th and 5fth also did the same. After the fifth one no. of breads remained are 3. Initially how many breads were there?

OpenStudy (anonymous):

x is number of breads

OpenStudy (anonymous):

total - breads taken=3

OpenStudy (anonymous):

i want to know that how much 1/2 of bread contributes to total number of breads

OpenStudy (calculusfunctions):

Alright @nitz I can teach you how to do this but you'll have listen (follow instructions) carefully and not interrupt me by typing while I am, and also answer my questions. Are you willing to do that and learn?

OpenStudy (anonymous):

here is the equation X-(5X/2+5/2X)=3

OpenStudy (anonymous):

solve it

OpenStudy (calculusfunctions):

Please stop giving out answers @Fazeelayaz and that's wrong anyway!

OpenStudy (calculusfunctions):

@nitz do you want to learn or not. If not then that's fine, there are plenty of other people I can teach. If yes then you need to reply to my post above.

OpenStudy (anonymous):

plz tell me

OpenStudy (anonymous):

total -breads taken=3 is that correct?

OpenStudy (calculusfunctions):

OK @Fazeelayaz but as I told @nitz earlier you'll have listen (follow instructions) carefully and not interrupt me by typing while I am, and also answer my questions. Are you willing to do that and learn? And no that is not correct! Now are you willing to learn or not?

OpenStudy (anonymous):

oh yes plz do it fast i am hungry

OpenStudy (calculusfunctions):

No! Proper learning and understanding cannot be rushed. Hence I'll ask one final time, do you or anyone else viewing this want to learn or not?!

OpenStudy (anonymous):

yes plz plz plz

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

plzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (calculusfunctions):

OK @Fazeelayaz and @nitz first as you correctly said at the beginning, let x represent the number of breads that were initially present. Now we shall approach this problem one step at a time. A step for each of the five thieves. Step 1: Thief #1 If there are initially x breads, and the 1st thief takes half of these plus half of one bread, then what is the expression that represents the number of breads that remain after this first action?

OpenStudy (anonymous):

X-(X/2+1/2X)

OpenStudy (anonymous):

x/2-1/2

OpenStudy (anonymous):

x-(1/2)x+(1/2)

OpenStudy (anonymous):

in brackets x/2-1/2

OpenStudy (calculusfunctions):

I'll get you started. Step 1: Thief #1 There are x breads, and if thief #1 takes half of these plus one-half of one, then the expression that represents the number of breads that remain after this first action is\[x -(\frac{ 1 }{ 2 }x +\frac{ 1 }{ 2 })=\frac{ 1 }{ 2 }x -\frac{ 1 }{ 2 }\]Understood?!

OpenStudy (calculusfunctions):

Very Good! @nitz

OpenStudy (anonymous):

now, what's next? as for the second: x/4-3/4

OpenStudy (calculusfunctions):

@irene22988 I'm trying to teach. Please stop giving out answers right now!

OpenStudy (anonymous):

O.o

OpenStudy (calculusfunctions):

I'm going to report anyone who gives out anymore answers, UNDERSTOOD?!

OpenStudy (anonymous):

i am just wondering: are we going to solve this by individually evaluating the number of remaining bread per thief?

OpenStudy (anonymous):

my bad. sorry.

OpenStudy (calculusfunctions):

Step 2: Thief #2 Now there are\[\frac{ 1 }{ 2 }x -\frac{ 1 }{ 2 }\]breads remaining. Thus thief #2 now starts with these many breads. If he takes half of these plus one-half of one bread, then he takes\[\frac{ 1 }{ 2 }(\frac{ 1 }{ 2 }x -\frac{ 1 }{ 2 })=\frac{ 1 }{ 4 }x +\frac{ 1 }{ 4 }\]Now the expression that represents the number of breads that remain after this second action is\[(\frac{ 1 }{ 2 }x -\frac{ 1 }{ 2 })-(\frac{ 1 }{ 4 }x +\frac{ 1 }{ 4 })=\frac{ 1 }{ 4 }x -\frac{ 3 }{ 4 }\]

OpenStudy (calculusfunctions):

Are we understanding thus far? If so then please humour me. Describe step 3: Thief #3, and thus tell me what is the expression that represents the number of breads that remain after this third action.

OpenStudy (anonymous):

Hahahahaha

OpenStudy (calculusfunctions):

I never said that @Algebraic! Why don't you wait 'til the end, and if I get it wrong (which guaranteed I won't), then you may get the satisfaction of criticizing me. Deal?

OpenStudy (calculusfunctions):

I'm not talking about a quarter of a bread, I'm talking about an expression.

OpenStudy (anonymous):

use the partial sum of a series.

OpenStudy (anonymous):

come on let him do what he is doing

OpenStudy (calculusfunctions):

@Algebraic! clearly they are not at that level, besides I didn't see you jumping in at the beginning. It's always very easy to be critical of others.

OpenStudy (calculusfunctions):

Thank you @Fazeelayaz

OpenStudy (calculusfunctions):

So do you know the expression that represents the number of breads that remain after this third action? That was my question to you before @Algebraic! rudely interrupted, while offering no real value.

OpenStudy (calculusfunctions):

@Fazeelayaz do you know the expression that represents the number of breads that remain after this third action or not?

OpenStudy (anonymous):

(x/4-3/4)-(x/8-3/8)=x/8-3/8

OpenStudy (calculusfunctions):

Almost but not quite! You inadvertently forgot to add one-half of a loaf to the expression for the number of loaves of bread the third thief takes. Care to give it another try @Fazeelayaz

OpenStudy (anonymous):

well i don't understand

OpenStudy (calculusfunctions):

That's alright! Allow me to explain. The second thief had left behind\[\frac{ 1 }{ 4 }x -\frac{ 3 }{ 4 }\]breads behind. Now the third thief takes half of these loaves of bread plus one-half of a loaf.Thus the third thief takes\[\frac{ 1 }{ 2 }(\frac{ 1 }{ 4 }x -\frac{ 3 }{ 4 })+\frac{ 1 }{ 2 }=\frac{ 1 }{ 8 }x +\frac{ 1 }{ 8 }\]Correct? Now the expression that represents the number of breads that remain after this third action is what? @Fazeelayaz ?

OpenStudy (anonymous):

x/4-3/4-(x/8+1/8)=x/8-7/8

OpenStudy (calculusfunctions):

Excellent!! @Fazeelayaz Do you now see what error you made earlier?

OpenStudy (anonymous):

yes

OpenStudy (calculusfunctions):

Great! @Fazeelayaz Now similarly can you tell me the expression that represents the number of loaves of bread that remain after the fourth action? Hence Step 4: Thief #4?

OpenStudy (anonymous):

x/16-7/16+1/2=x/16+1/16 x/8+1/8-(x/16+1/16)=x/16+1/16

OpenStudy (calculusfunctions):

@Fazeelayaz that's the expression for the number of loaves of bread that thief #4 takes. What is the expression that represents the number of loaves of bread that remain after this fourth action?

OpenStudy (anonymous):

same

OpenStudy (calculusfunctions):

No @Fazeelayaz please try again.

OpenStudy (anonymous):

x/16-7/16+1/2=x/16+1/16 x/8-7/8-(x/16+1/16)=x/16-15/16

OpenStudy (calculusfunctions):

Correct!! Very Good! Now can you tell me the expression that represents the number of loaves of bread that remain after the fifth action?

OpenStudy (anonymous):

x/32-15/32+1/2=x/32+1/32 x/16-15/16-(x/32+1/32)=x/32-31/32

OpenStudy (calculusfunctions):

@Fazeelayaz YES!! Now after this fifth action, there were 3 breads left, thus equate this final expression to 3, and solve for x. Go ahead @Fazeelayaz

OpenStudy (anonymous):

x/32-31/32=3 x-31=96 x=96+31 x=127

OpenStudy (calculusfunctions):

PERFECT!!! SEE, YES YOU CAN!! EXCELLENT!!! ∴ there were initially 127 loaves of bread!

OpenStudy (anonymous):

thank u very much

OpenStudy (anonymous):

o thanks for that too

OpenStudy (anonymous):

well i have to go iam extremelly hungry THANKS FOR UR TIME SIR

OpenStudy (anonymous):

Im glad dat @Fazeelayaz got it... Cheer up Hip hip HOOORAY =D =D

OpenStudy (calculusfunctions):

You're very welcome, we could have finished this way before if others hadn't been obnoxious and kept rudely interrupting but that's okay. What matters most is that you @Fazeelayaz , @nitz , and @Algebraic! learned how to properly solve these types of problems. Excellent work @Fazeelayaz !!!!!!!

OpenStudy (calculusfunctions):

@Fazeelayaz is there anything else I can help you with?

OpenStudy (anonymous):

Well done @calculusfunctions u taught alot of people..SHOWERS OF BLESSINGS =D

OpenStudy (calculusfunctions):

Thank you @Ivanka123

OpenStudy (anonymous):

Ur most welcome =D